In an explosive volcanic eruption, a molten rock spewed _ out from the craterof the volcano reaches a maximum height of 2.481 km above ground. After 37.0 since it WaS released, it hit the ground, distance 5.25 km away from the crater: How much time did it took the molten rock "iio t0 reach the maximum height? () What is the magnitude and direction of the velocity of the molten rock as it was Ay thrown out from the crater? (c) How high is the volcano? 5.25 km (d) What is the magnitude and direction of the velocity of the molten rock when it landed on the ground?

1
Expert's answer
2022-03-07T10:19:22-0500

a)

"t=15.4~s,"

b)

"v=70.7~\frac ms,"

"\alpha=24°,"

c)

"h=3.06~km,"

d)

"u=101.2~\frac ms,"

"\beta=70.2°."